# TransitReID: Transit OD Data Collection with Occlusion-Resistant Dynamic Passenger Re-Identification

### v1

@@ Abstract
Transit Origin-Destination (OD) data are essential for transit planning, particularly in route optimization and demand-responsive paratransit systems. Traditional methods, such as manual surveys, are costly and inefficient, while Bluetooth and WiFi-based approaches require passengers to carry specific devices, limiting data coverage. On the other hand, most transit vehicles are equipped with onboard cameras for surveillance, offering an opportunity to repurpose them for edge-based OD data collection through visual person re-identification (ReID). However, such approaches face significant challenges, including severe occlusion and viewpoint variations in transit environments, which greatly reduce matching accuracy and hinder their adoption. Moreover, designing effective algorithms that can operate efficiently on edge devices remains an open challenge. To address these challenges, we propose TransitReID, a novel framework for individual-level transit OD data collection. TransitReID consists of two key components: (1) An occlusion-robust ReID algorithm featuring a variational autoencoder guided region-attention mechanism that adaptively focuses on visible body regions through reconstruction loss-optimized weight allocation; and (2) a Hierarchical Storage and Dynamic Matching (HSDM) mechanism specifically designed for efficient and robust transit OD matching which balances storage, speed, and accuracy. Additionally, a multi-threaded design supports near real-time operation on edge devices, which also ensuring privacy protection. We also introduce a ReID dataset tailored for complex bus environments to address the lack of relevant training data. Experimental results demonstrate that TransitReID achieves state-of-the-art performance in ReID tasks, with an accuracy of approximately 90% in bus route simulations.
@@ I Introduction
Transit Origin-Destination (OD) data collection plays a critical role in the fields of transportation engineering for urban planning, service frequency adjustments, and route optimization. It aids in the analysis of relationships and impacts among regional economic development, infrastructure, and urban mobility [1]. This data represents the demand of passengers and is needed to design the bus routes efficiently to minimize travel time, reduce operating costs, and driver scheduling. As shown in Fig. 1, traditional methods for estimating transit OD data rely on manual approaches such as surveys, which are labor-intensive and suffer from low response rates [2].
More advanced methods, such as those utilizing mobile phone data[3] and Bluetooth technology [4], require passengers to carry specific devices with WiFi or Bluetooth enabled, which limits coverage and might raise privacy concerns due to the unique identifiers of the device. Automated Passenger Counters (APCs) [5] also assist in estimating OD data, however, they can only capture the aggregate number and locations of boarding and alighting passengers. They fail to match passengers at the individual level, which is crucial for studies focused on individual passenger behavior.
Meanwhile, most transit vehicles in the US are already equipped with onboard cameras for liability and surveillance purposes, presenting a valuable opportunity to repurpose them as smart sensors for transit OD data collection. By integrating these cameras with person re-identification (ReID) techniques and edge computing devices, seamless individual-level matching can be achieved [6]. By integrating sensors such as GPS, the matching results can be readily translated into transit OD data.
Despite the advantages of camera-based methods over prevailing techniques, their adoption in transit OD data collection remains limited due to challenges posed by the transit vehicle environment, particularly occlusion issues. For example, during peak times, boarding and alighting passengers may obscure each other from the camera view, leading to information loss and background confusion. Recent research increasingly focuses on occlusion issues in ReID, broadly categorized into two types: methods based on local attention mechanisms and those based on feature recovery. The former constructs attention mechanisms by focusing on different local areas to suppress interference from occluded parts [7], while the latter attempts to recover information from occluded areas by associating temporal or spatial information [8, 9]. Fixed local regions are used to differentiate various body parts of pedestrians and assign scores to different areas using a quality predictor in [10]. However, this fixed-area approach is not suitable for passenger matching tasks on transit vehicles under dynamic viewpoints. As the distance and angles between passengers and cameras change over time, the regions of their body parts shift throughout the video.
@@
Due to limitations in power consumption, memory, and computational resources, edge devices are generally incapable of executing advanced and complex analytical algorithms[11, 12]. Fast and accurate estimation of OD (origin-destination) data is essential for enabling real-time passenger flow analysis in transit systems. However, most deep learning algorithms have relatively long inference times on edge devices, leading to latency and performance degradation. Some efforts have focused on model lightweighting techniques, such as knowledge distillation and pruning, or on optimizing neural network architectures specifically for edge hardware[13]. Nevertheless, lightweight models often suffer from a trade-off in performance, with reduced accuracy to some extent.
@@
What’s more, unlike traditional ReID studies, our work focuses on the problem of transit OD data collection, hence the gallery, shown in Fig. 2, is dynamic, changing with the variations in passengers boarding and alighting at each station. Consequently, ReID evolves from a problem of matching 1 query to $N$ options, into a dynamic 1 to $N_{1},N_{2},...N_{K}$ problem, where $N_{k}$ represents the size of the gallery updated at station $k$. To address this, we design a hierarchical storage and dynamic matching (HSDM) method to balance the trade-offs among matching accuracy, storage pressure, and processing speed.
@@
To jointly tackle the aforementioned issues of occlusion and the constraints of online adaptive matching, we propose TransitReID, an individual-level and robust transit OD data collection method that incorporates occlusion-resistant and dynamic passenger ReID. The overall framework is shown in Fig. 3, featuring two principal technical innovations: (1) An occlusion-robust ReID algorithm that employs a region-attention module to adaptively weight less obscured body segments with clearer viewpoints, guided by a co-trained Variational Autoencoder (VAE) grading model. The VAE’s reconstruction loss directly informs attention weight allocation, emphasizing visually coherent regions. This is followed by a feature fusion strategy that preserves discriminative information through concatenated weighted averages of region embeddings. (2) A Hierarchical Storage and Dynamic Matching (HSDM) mechanism specifically designed for efficient transit OD matching, which incorporates Hot/Cold Storage and a Snatch mechanism to balance accuracy, storage, and speed. A multi-threaded design is implemented to support near real-time operation on transit vehicles equipped with edge devices, without requiring any modifications to the models or architectures. Notably, the framework processes video locally on edge devices and stores only high-level, non-identifiable features, ensuring a privacy-preserving and non-intrusive solution.
@@
Another issue is that current video ReID datasets [14, 15, 16, 17] are not specifically designed for occlusion scenarios, while most occluded person ReID datasets [18, 19, 20, 21] are image-based and not suited for training tasks under transit surveillance video conditions. Additionally, person images in those datasets are captured from consistent angles, whereas in transit scenarios, the angles of passenger images captured by surveillance cameras change continuously. This variability is crucial for training a robust and generalizable model under transit bus operational scenarios. Therefore, we construct a passenger ReID dataset collected from real bus operational settings, including footage from cameras at the front and rear doors. This dataset captures 17,636 images of 157 passengers.
In essence, the major contributions of this work are as follows:
  • Onboard transit surveillance systems are combined with computer vision technology to collect transit OD data, achieving individual-level precision in the construction of the OD matrix.
  • An attention-based passenger ReID algorithm is specifically designed for transit vehicles scenarios with variable occlusion and viewpoints.
@@
  • An efficient programming approach is designed to enable real-time operation on edge devices for smart transit applications. In addition, edge-based processing inherently supports privacy protection.
Figure 2: Segmentation results after detection and tracking.
@@ II-A Transit OD Data Collection
Transit OD data collection refers to the systematic process of gathering information about the starting points (origins) and endpoints (destinations) of individual trips within a public transportation network. The most traditional and ubiquitous methods for estimating transit OD data is by surveys [22, 23], which are labor-intensive and have limited response rates. In [24], an integrated program based on Automated Fare Collection (AFC) systems was developed that can generate OD matrices and passenger load profiles by making some assumptions about round trips. A trajectory search algorithm was proposed to process passenger smart card data to track daily travel trajectories in [25]. An optimization model was proposed in [26] to approximate network-level OD flows with big data generated from Automated Passenger Counters (APCs). RFID technology was utilized to create a Boarding-In and Boarding-Out (BIBO) system for passenger monitoring in [27].
Moreover, mobile phone data [3], Bluetooth technology [4], and multi-source approaches – such as combining GIS data with passenger IC card data to construct OD matrices, as demonstrated in [28] – have been utilized. However, these methods face challenges, including high costs, significant maintenance requirements [2], and the need for passengers to carry specific devices, which limits data coverage. Additionally, the extensive use of these systems in public transport can lead to privacy concerns [29], such as exposing details about ticket types, passenger categories, travel behaviors, and purchasing methods.
While most methods capture only the number and locations of boarding and alighting passengers, they lack the precision to individually track and match passengers, crucial for detailed behavioral studies. Video-based passenger re-identification (ReID) technology combined with transit surveillance cameras provides a potential solution to this problem. In [6], the YOLO algorithm is utilized to detect passengers, generating feature vectors through a neural network and selecting matches based on vector distance, but this work ignores the passive influence of occlusion. Several works [30, 31, 32] address the problem by only collecting information from head images of passengers inside the bus. However, this approach overlooks the potential rich information from other regions, lacks discrimination power against similar headgear, and is not robust to changes in lighting.
@@ II-B Occluded Re-Identification
Occlusion is by far the major problem in the computer vision-based person ReID task. Most of the occluded person re-identification methods rely on convolutional neural networks (CNNs)[7]. These aim at body part-level feature measurement, for example, a region-based quality predictor was proposed and trained with the regional feature generation model to get the weighted feature representation in [10]. However, the regions defined in this work is fixed, thus not suitable for dynamic perspectives. A pose-guided feature alignment (PGFA)[18] utilized human pose landmarks to build an attention map, guiding the model to extract useful information from non-occluded parts. Similarly, [33, 34, 35, 36] utilize posture-guided attention mechanisms to discern visible areas. These methods partition the ReID problem into two segments: initially detecting key areas, then selectively extracting regional features for matching. However, the accuracy of the posture estimation model can directly impact subsequent discrimination. Other methods include feature recovery[8, 9, 37, 38, 39], graph convolution[40], and multi-model methods[41, 42]. Unfortunately, these methods often struggle with generalization across diverse scenarios, such as changes in viewpoints, variations in environmental lighting, and complex occlusions among passengers in transit vehicle environments. Additionally, the lack of scenario-specific training data further limits the models’ generalizability. Moreover, they address static ReID problems, whereas the gallery in a transit vehicle operating environment is inherently dynamic. We design a hierarchical storage and dynamic matching (HSDM) mechanism specifically adapted for dynamic ReID, detailed in Section III. F.
Figure 3: The Framework of the Proposed TransitReID. In Dynamic Transit ReID, the surveillance system captures image sequences of the same passenger boarding and alighting at different stops. These images are processed by a feature generation network to extract one-dimensional feature representations. If the passenger is boarding, the features are temporarily stored in the gallery. If the passenger is alighting, the system retrieves a matching result from the gallery through the hierarchical storage and dynamic matching (HSDM) mechanism. In the Feature Generation Network, the input consists of an image sequence of a passenger. After action detection and segmentation, each mask is processed by a human parsing network, which divides the image into three body parts: ‘head’, ‘torso’, and ‘legs’. Each part is then fed into a grading model and a baseline model to obtain a quality score and an initial feature representation. Finally, a feature averaging unit integrates these outputs into the final 3*2048 length feature vector.
@@ III-A Problem Definition
In the following sections, we focus on the transit bus environment, which is equipped with two cameras: a front-door camera $C_{f}$ and a rear-door camera $C_{r}$, which capture sequences of images over a time interval $T$. The image sets captured by each camera are:
$I_{f}=\{i_{f,1},i_{f,2},...,i_{f,T}\},\quad I_{r}=\{i_{r,1},i_{r,2},...,i_{r,T}\}.$
Each image may contain multiple passengers. The gallery, $G$ consists of all passenger images extracted from the raw images recorded by both cameras while onboard the bus:
$G=G_{f}\cup G_{r}=\{g_{f,1},g_{f,2},...,g_{f,M_{f}}\}\cup\{g_{r,1},g_{r,2},...,g_{r,M_{r}}\},$
@@
At each bus stop, the query set, $Q$, consists of passengers departing from either the front or rear door:
$Q=Q_{f}\cup Q_{r}=\{q_{f,1},q_{f,2},...,q_{f,N_{f}}\}\cup\{q_{r,1},q_{r,2},...,q_{r,N_{r}}\},$
where $N_{f}$ and $N_{r}$ are the number of passengers alighting from the front and rear doors, respectively, and $N=N_{f}+N_{r}$ is the total number of alighting passengers.
Each passenger $p_{k}$ is mapped to a high-dimensional feature space via a function:
$\phi:p_{k}\rightarrow\mathbb{R}^{d},$
where $d$ is the feature space dimension. For each alighting passenger $q_{k}\in Q$, we determine the best match from the gallery $G$ by computing the distance between $\phi(q_{k})$ and $\phi(g_{j})$ for all $g_{j}\in G$, selecting the closest match:
$\mathbf{min}\ distance(\phi(q_{k}),\phi(g_{j})),\quad\forall g_{j}\in G.$

### v2

@@ Abstract
Transit Origin-Destination (OD) data are fundamental for optimizing public transit services, yet current collection methods, such as manual surveys, Bluetooth/WiFi tracking, or Automated Passenger Counters, are either costly, device-dependent, or incapable of individual-level matching. Meanwhile, onboard surveillance cameras already deployed on most transit vehicles provide an underutilized opportunity for automated OD data collection. Leveraging this, we present TransitReID, a novel framework for individual-level and occlusion-resistant passenger re-identification (ReID) tailored to transit environments. Our approach introduces three key innovations: (1) an occlusion-robust ReID algorithm that integrates a variational autoencoder–guided region-attention mechanism and selective quality feature averaging to dynamically emphasize visible and discriminative body regions under severe occlusions and viewpoint variations; (2) a Hierarchical Storage and Dynamic Matching (HSDM) mechanism that transforms static gallery matching into a dynamic process, balancing accuracy, memory, and speed in real-world bus operations; and (3) a multi-threaded edge implementation that enables near real-time OD estimation while ensuring privacy by processing all data locally. To support research in this domain, we also construct a new Transit ReID dataset with over 17,000 images captured from bus front/rear cameras under diverse occlusion and viewpoint conditions. Experimental results demonstrate that TransitReID achieves state-of-the-art performance, with R-1 accuracy of 88.3% and mAP of 92.5%, and further sustains 90% OD estimation accuracy in bus route simulations on NVIDIA Jetson edge devices. This work advances both the algorithmic and system-level foundations of automated transit OD collection, paving the way for scalable, privacy-preserving deployment in intelligent transportation systems.
@@ I Introduction
Transit Origin-Destination (OD) data collection plays a critical role in the fields of transportation engineering for urban planning, service frequency adjustments, and route optimization. It aids in the analysis of relationships and impacts among regional economic development, infrastructure, and urban mobility [1]. For public transportation systems, OD data represents passenger demand and plays a crucial role in designing routes more efficiently to minimize travel time, reduce operating costs, and optimize driver scheduling.
As shown in Fig. 1, traditional methods for estimating transit OD data rely on manual approaches such as surveys, which are labor-intensive and suffer from low response rates [2]. More advanced methods, such as those utilizing mobile phone data[3] and Bluetooth technology [4], require passengers to carry specific devices with WiFi or Bluetooth enabled, which limits coverage and might raise privacy concerns due to the unique identifiers of the device. Automated Passenger Counters (APCs) [5] also assist in estimating OD data, however, they can only capture the counts and locations of boarding and alighting passengers and fail to match passengers at the individual level, which is crucial for understanding detailed travel behavior.
From a transit planner’s perspective, individual-level OD information provides significant advantages over aggregated OD data. It enables more accurate demand forecasting by revealing detailed movement patterns, supports the design of tailored services for different passenger groups such as commuters, students, and occasional travelers, and facilitates the evaluation of OD-related performance indicators across user segments. Moreover, it helps identify irregular behaviors, such as fare evasion or abnormal travel trajectories, and can assist law enforcement in locating suspects or missing persons, thereby improving both operational efficiency and public safety.
@@
To address these needs, video-based re-identification (ReID) techniques offer a promising solution. Unlike device-dependent approaches, video-based ReID captures continuous passenger trajectories without requiring user cooperation and, when combined with anonymization strategies, can achieve this while mitigating privacy concerns. Integrating ReID into OD estimation frameworks can thus significantly enhance the accuracy, coverage, and utility of OD data, enabling transit planners to design smarter and more responsive public transportation systems. Meanwhile, most transit vehicles in the US are already equipped with onboard cameras for liability and surveillance purposes, presenting a valuable opportunity to repurpose them as smart sensors for video-based transit OD data collection [6]. By integrating these cameras with ReID techniques, edge computing devices, and GPS, seamless individual-level matching and OD data collection can be achieved.
Despite the advantages of camera-based methods over existing techniques, their adoption in transit OD data collection remains limited due to several challenges, particularly occlusion. For example, during peak times, passengers may obscure each other from the camera view, leading to information loss and background confusion. Recent research on ReID has increasingly emphasized occlusion, with existing approaches broadly falling into two categories: local attention–based methods and feature recovery–based methods. The former constructs attention mechanisms by focusing on different local areas to suppress interference from occluded parts [7], while the latter attempts to recover information from occluded areas by associating temporal or spatial information [8, 9]. Although some methods achieve strong performance on general ReID benchmarks, they still struggle with dynamic viewpoints and complex occlusion patterns, and often fail to transfer effectively to real-world application scenarios such as transit vehicles.
@@
Additionally, unlike traditional ReID studies, our work focuses on the problem of transit OD data collection, hence the gallery, shown in Fig. 2, is dynamic, changing with the variations in passengers boarding and alighting at each station. Consequently, transit ReID evolves from a problem of matching 1 query to $N$ options, into a dynamic 1 to $(N_{1},N_{2},...,N_{K})$ problem, where $N_{k}$ represents the size of the gallery updated at station $k\in K$. To address this, we design a hierarchical storage and dynamic matching (HSDM) method to balance the trade-offs among matching accuracy, storage pressure, and processing speed.
The limited computational resources of edge devices present an additional challenge in executing advanced and complex analytical algorithms [10, 11]. Fast and accurate estimation of OD data is crucial for minimizing latency in passenger flow analysis. To address this, some efforts have focused on model lightweighting techniques, such as knowledge distillation and pruning, or on optimizing neural network architectures specifically for edge hardware [12]. Nevertheless, lightweight models often face a trade-off between efficiency and accuracy, typically leading to a certain degree of performance degradation.
Figure 2: Segmentation results after detection and tracking.
@@
To jointly tackle the aforementioned issues, we propose TransitReID, an individual-level and robust transit OD data collection method that incorporates occlusion-resistant and dynamic passenger ReID. The overall framework is shown in Fig. 3, featuring two principal technical innovations: (1) An occlusion-robust ReID algorithm employing a region-attention module to adaptively weight less-obscured body segments with clearer viewpoints, guided by a co-trained Variational Autoencoder (VAE) grading model robust to occlusion and viewpoint variations. The VAE’s reconstruction loss directly informs attention weight allocation, emphasizing visually coherent regions. This is followed by a Selective Quality Feature Averaging (SQFA) strategy that preserves discriminative information through concatenated weighted averages of region embeddings. (2) A Hierarchical Storage and Dynamic Matching (HSDM) mechanism specifically designed for efficient transit OD matching, which incorporates Hot/Cold Storage and a Snatch mechanism to balance accuracy, storage, and speed. A multi-threaded design is implemented to support near real-time operation on transit vehicles equipped with edge devices, without requiring any modifications to the models or architectures. Notably, the framework processes video locally on edge devices and stores only high-level, non-identifiable features, ensuring a privacy-preserving and non-intrusive solution.
@@
Another issue is that current video ReID datasets [13, 14, 15, 16] are not specifically designed for occlusion scenarios, while most occluded person ReID datasets [17, 18, 19, 20] are image-based and not suited for training tasks under transit surveillance video conditions. Additionally, person images in those datasets are captured from consistent angles, whereas in transit scenarios, the angles of passenger images captured by surveillance cameras change continuously. This variability is crucial for training a robust and generalizable model under transit bus operational scenarios. Therefore, we construct a transit ReID dataset (detailed in Section IV. A.) collected from real bus operational settings, including 17,636 images of 157 passengers captured by cameras at the front and rear doors.
The major contributions of this work are as follows:
  • We propose TransitReID to enable individual-level transit OD data collection, employing an attention-based passenger ReID algorithm that is robust to variable occlusions and viewpoint changes.
@@
  • Introduce an efficient programming approach to enable real-time operation on edge devices for smart transit applications. Edge-based processing inherently supports privacy protection.
Figure 3: The framework of the proposed TransitReID. The surveillance system captures image sequences of the passengers boarding and alighting at different stops. After behavior detection, each image is processed by a human parsing network, which divides the image into three body parts: ‘head’, ‘torso’, and ‘legs’. Each part is then fed into a grading model and a feature generation model to obtain a quality score and an initial feature representation. Finally, a selective quality feature averaging (SQFA) unit integrates these outputs into the final feature vector. At boarding, the feature is first stored in the gallery; at alighting, the system employs the hierarchical storage and dynamic matching (HSDM) mechanism to identify the corresponding feature from the gallery.
@@ II-A Transit OD Data Collection
Transit OD data collection refers to the systematic process of gathering information on the origins and destinations of individual trips within a public transportation network. OD trip matrices can be estimated either from direct trip observations or derived from traffic counts. The latter approach, known as OD synthesis, estimates the OD matrix through linear combinations of observed flows [21]. Common methods include gravity-based models [22], entropy models [23, 24], and statistical approaches based on least squares [25] or maximum likelihood estimation [26]. However, these methods are typically applied to vehicular traffic.
The most ubiquitous methods for estimating passenger-level transit OD data is by surveys [27, 28], which are labor-intensive and have limited response rates. In [29], an integrated program based on Automated Fare Collection (AFC) systems was developed that can generate OD matrices and passenger load profiles by making some assumptions about round trips. A trajectory search algorithm was proposed to process passenger smart card data to track daily travel trajectories in [30]. An optimization model was proposed in [31] to approximate network-level OD flows with big data generated from Automated Passenger Counters (APCs). RFID technology was utilized to create a Boarding-In and Boarding-Out (BIBO) system for passenger monitoring in [32].
Moreover, mobile phone data [3], Bluetooth technology [4], and multi-source approaches — such as integrating GIS data with passenger IC card data to construct OD matrices, as demonstrated in [33] — have been employed. However, these methods face several challenges, including high costs, substantial maintenance requirements [2], and the dependence on passengers carrying specific devices, which limits data coverage. In addition, widespread deployment of such systems in public transportation raises privacy concerns [34], potentially revealing sensitive information.
While most existing methods capture only aggregate counts and locations, they lack the precision required to track and match individual passengers, which is essential for detailed behavioral analysis. Video-based passenger re-identification using transit surveillance cameras offers a promising solution. In [6], the YOLO algorithm was employed to detect passengers, followed by feature extraction through a neural network and matching based on vector distances. However, this study does not consider the passive effects of occlusion. Other works [35, 36, 37] addressed this issue by focusing solely on head images of passengers inside the bus. Yet, they neglects potentially informative regions of the body, exhibits limited discriminative power against similar headgear, and is sensitive to variations in lighting conditions.
@@ II-B Occluded Re-Identification
Occlusion is one of the most significant challenges in vision-based person ReID. Some of the occluded person re-identification methods rely on convolutional neural networks (CNNs)[7]. They aim at body part-level feature measurement, for example, a region-based quality predictor was proposed and trained with the regional feature generation model to get the weighted feature representation in [38]. However, the regions defined in this work is fixed, thus not suitable for passenger matching tasks on transit vehicles under dynamic viewpoints. As the distance and angles between passengers and cameras change over time, the regions of their body parts shift throughout the video. A pose-guided feature alignment method (PGFA)[17] utilized human pose landmarks to build an attention map, guiding the model to extract useful information from non-occluded parts. Similarly, [39, 40, 41, 42] utilized posture-guided attention mechanisms to discern visible areas. These methods partition the ReID problem into two segments: initially detecting key areas, then selectively extracting regional features for matching. However, the accuracy of the posture estimation model can directly impact subsequent discrimination. Other methods include feature recovery[8, 9, 43, 44, 45], graph convolution[46], and multi-model methods[47, 48]. Unfortunately, these methods often struggle with generalization across diverse scenarios, such as changes in viewpoints, variations in environmental lighting, and complex occlusions among passengers in transit vehicle environments. Additionally, the lack of scenario-specific training data further limits the models’ generalizability. Moreover, they address static ReID problems, whereas the gallery in a transit vehicle is inherently dynamic. We design a hierarchical storage and dynamic matching (HSDM) mechanism specifically adapted for dynamic ReID, detailed in Section III. F.
@@ III-A Problem Definition
In the following sections, we focus on the transit bus environment, which is equipped with two cameras: a front-door camera $C_{f}$ and a rear-door camera $C_{r}$, which capture sequences of images over a time interval $T$. Each image may contain multiple passengers. The gallery $G$ consists of all passenger images extracted from the raw images recorded by both cameras while onboard the bus:
$G=G_{f}\cup G_{r}=\{g_{f,1},g_{f,2},...,g_{f,M_{f}}\}\cup\{g_{r,1},g_{r,2},...,g_{r,M_{r}}\}$ (1)
@@
At each bus stop, the query $Q$ updates passengers departing from either the front or rear door:
$Q=Q_{f}\cup Q_{r}=\{q_{f,1},q_{f,2},...,q_{f,N_{f}}\}\cup\{q_{r,1},q_{r,2},...,q_{r,N_{r}}\}$ (2)
where $N_{f}$ and $N_{r}$ are the number of passengers alighting from the front and rear doors, respectively, and $N=N_{f}+N_{r}$ is the total query size.
Each passenger $p$ is mapped to a high-dimensional feature space via a function:
$\phi:p\rightarrow\mathbb{R}^{d}$ (3)
where $d$ is the feature space dimension. For each alighting passenger $q_{n}\in Q$, we determine the best match from the gallery $G$ by computing the distance between $\phi(q_{n})$ and $\phi(g_{m})$ for all $g_{m}\in G$, selecting the closest match:
$\hat{g}=\arg\min_{g_{m}\in G}\,distance\big{(}\phi(q_{n}),\phi(g_{m})\big{)}$ (4)